1. Suppose X has mean 3 and variance 4, Y has mean 5 and variance 9, and Coy(X, Y) =-2. (a) What is the mean of 6X 7Y? (b) What is the variance of 6X TY? (c) What is the variance of 6X - TY? (d) What is the squared coefficient of variation of X? (e) What is the covariance of X and X +Y?
Answer #1

a)

-(6x3)+(7x5) = 18+35

b)

36V (x) 49V() +84cov(X, Y) 36 × 4+49 × 9-168 = 144 + 441-168 417

c)

36V (x) 49V (X) - 84cov(X, Y) = 36 × 4 +49 × 9 + 168 = 144 + 441 + 168 = 753

d)

The squared coefficient of variation is calculated below:

NE 2 2-34-9